# Inexact Moreau Envelope Lagrangian Method for Non-Convex Constrained Optimization under Local Error Bound Conditions on Constraint Functions

### v1

@@ 5 Convergence Analysis for the iMELa Method
$\tau_{t}=\tau=\frac{p-L}{4B_{g}^{2}}\text{ and }\theta_{t}=\theta\leq\frac{p-L}{18p},\leavevmode\nobreak\ \forall\,t\geq 0.$
@@
$\phi^{t}-\phi^{t+1}\geq$ $\leavevmode\nobreak\ \frac{p-L}{4}\|{\mathbf{x}}^{(t)}-\tilde{{\mathbf{x}}}^{(t+1)}\|^{2}-\frac{\epsilon_{t}^{2}}{2(p-L)}-6p\theta\|\tilde{{\mathbf{x}}}^{(t+1)}-{\mathbf{x}}({\mathbf{z}}^{(t)})\|^{2}$
@@
$+\frac{p}{6\theta}\|{\mathbf{z}}^{(t)}-{\mathbf{z}}^{(t+1)}\|^{2}+\tau\cdot\textup{dist}^{2}\big{(}{\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)}),\mathcal{N}_{\mathbb{R}_{+}^{m}}(\boldsymbol{\lambda}^{(t+1)})\big{)}.$ (41)
Proof.
Applying (35), (39) and (40) to (33) and rearranging terms, we have
$\phi^{t}-\phi^{t+1}=$ $\leavevmode\nobreak\ \mathcal{L}_{p}({\mathbf{x}}^{(t)},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t)})-\mathcal{L}_{p}({\mathbf{x}}^{(t+1)},{\mathbf{z}}^{(t+1)},\boldsymbol{\lambda}^{(t+1)})$
$\geq$ $\leavevmode\nobreak\ \frac{p-L}{2}\|{\mathbf{x}}^{(t)}-\tilde{{\mathbf{x}}}^{(t+1)}\|^{2}-\frac{\epsilon_{t}^{2}}{2(p-L)}+\frac{p}{2}(2/\theta_{t}-1)\|{\mathbf{z}}^{(t)}-{\mathbf{z}}^{(t+1)}\|^{2}$
$-p(1/\sigma+1)\|{\mathbf{z}}^{(t)}-{\mathbf{z}}^{(t+1)}\|^{2}.$ (42)
For any $\alpha>0$, by Young’s inequality, it holds that
By Cauchy-Schwartz inequality and (32), we have
$\leavevmode\nobreak\ 2\langle{\mathbf{z}}^{(t+1)}-{\mathbf{z}}^{(t)},{\mathbf{x}}(\boldsymbol{\lambda}^{(t+1)},{\mathbf{z}}^{(t+1)})-\tilde{{\mathbf{x}}}^{(t+1)}\rangle$
$\leq$ $\leavevmode\nobreak\ 2\|{\mathbf{z}}^{(t+1)}-{\mathbf{z}}^{(t)}\|\cdot\|{\mathbf{x}}(\boldsymbol{\lambda}^{(t+1)},{\mathbf{z}}^{(t+1)})-\tilde{{\mathbf{x}}}^{(t+1)}\|\leq(2/\sigma)\cdot\|{\mathbf{z}}^{(t+1)}-{\mathbf{z}}^{(t)}\|^{2}.$ (44)
Applying (43) and (44) to (42), we have
$\phi^{t}-\phi^{t+1}\geq$ $\leavevmode\nobreak\ \frac{p-L}{2}\|{\mathbf{x}}^{(t)}-\tilde{{\mathbf{x}}}^{(t+1)}\|^{2}-\frac{\epsilon_{t}^{2}}{2(p-L)}-p\alpha\|\tilde{{\mathbf{x}}}^{(t+1)}-{\mathbf{x}}({\mathbf{z}}^{(t)})\|^{2}$
Therefore, we have
$\leavevmode\nobreak\ 2\langle\boldsymbol{\lambda}^{(t+1)}-\boldsymbol{\lambda}^{(t)},{\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)})-{\mathbf{g}}({\mathbf{x}}^{(t)})\rangle+\langle\boldsymbol{\lambda}^{(t+1)}-\boldsymbol{\lambda}^{(t)},{\mathbf{g}}({\mathbf{x}}^{(t)})\rangle$
$\geq$ $\leavevmode\nobreak\ 2\langle\boldsymbol{\lambda}^{(t+1)}-\boldsymbol{\lambda}^{(t)},{\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)})-{\mathbf{g}}({\mathbf{x}}^{(t)})\rangle+\tau_{t}^{-1}\cdot\|\boldsymbol{\lambda}^{(t+1)}-\boldsymbol{\lambda}^{(t)}\|^{2}$
$=$ $-\tau_{t}\|{\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)})-{\mathbf{g}}({\mathbf{x}}^{(t)})\|^{2}+\tau_{t}^{-1}\|\boldsymbol{\lambda}^{(t+1)}-\boldsymbol{\lambda}^{(t)}+\tau_{t}({\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)})-{\mathbf{g}}({\mathbf{x}}^{(t)}))\|^{2}$
$\geq$ $-\tau_{t}\|{\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)})-{\mathbf{g}}({\mathbf{x}}^{(t)})\|^{2}+\tau_{t}^{-1}\cdot\text{dist}^{2}\big{(}\tau_{t}{\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)}),\mathcal{N}_{\mathbb{R}_{+}^{m}}(\boldsymbol{\lambda}^{(t+1)})\big{)}$
$=$ $-\tau_{t}\|{\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)})-{\mathbf{g}}({\mathbf{x}}^{(t)})\|^{2}+\tau_{t}\cdot\text{dist}^{2}\big{(}{\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)}),\mathcal{N}_{\mathbb{R}_{+}^{m}}(\boldsymbol{\lambda}^{(t+1)})\big{)},$ (48)
Applying (48) and Assumption 3.1E to (45), we have
$+\tau_{t}\cdot\text{dist}^{2}\big{(}{\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)}),\mathcal{N}_{\mathbb{R}_{+}^{m}}(\boldsymbol{\lambda}^{(t+1)})\big{)}.$ (49)
According to Proposition 5.5, we notice that to ensure a sufficient decrease of $\phi_{t}$, we need to further eliminate the negative term $-6p\theta\|\tilde{{\mathbf{x}}}^{(t+1)}-{\mathbf{x}}({\mathbf{z}}^{(t)})\|^{2}$ in (41). To do so, we will employ the local error bound condition assumed in Assumption 3.3.
Lemma 5.6.
$\textup{dist}^{2}\big{(}-\nabla f(\tilde{{\mathbf{x}}})-\nabla{\mathbf{g}}(\tilde{{\mathbf{x}}})\boldsymbol{\lambda},\mathcal{N}_{\mathcal{X}}(\tilde{{\mathbf{x}}})\big{)}+\textup{dist}^{2}\big{(}{\mathbf{g}}(\tilde{{\mathbf{x}}}),\mathcal{N}_{\mathbb{R}_{+}^{m}}(\boldsymbol{\lambda})\big{)}\leq R^{2}(\delta^{\prime})$ (51)
Proof.
$\text{dist}^{2}\big{(}-\nabla f(\tilde{{\mathbf{x}}}^{(k)})-\nabla{\mathbf{g}}(\tilde{{\mathbf{x}}}^{(k)})\tilde{\boldsymbol{\lambda}}^{(k)},\mathcal{N}_{\mathcal{X}}(\tilde{{\mathbf{x}}}^{(k)})\big{)}+\text{dist}^{2}\big{(}{\mathbf{g}}(\tilde{{\mathbf{x}}}^{(k)}),\mathcal{N}_{\mathbb{R}_{+}^{m}}(\tilde{\boldsymbol{\lambda}}^{(k)})\big{)}\leq 1/k,$ (52)
Due to the compactness of $\mathcal{X}$ (see Assumption 3.1A) and the set $\{\boldsymbol{\lambda}\in\mathbb{R}_{+}^{m}:\|\boldsymbol{\lambda}\|\leq M_{\boldsymbol{\lambda}}\}$, by passing to a subsequence if necessary, there exists some pair $(\tilde{{\mathbf{x}}}^{*},\tilde{\boldsymbol{\lambda}}^{*})\in\mathcal{X}\times\mathbb{R}_{+}^{m}$, $\|\tilde{\boldsymbol{\lambda}}^{*}\|\leq M_{\boldsymbol{\lambda}}$, such that $(\tilde{{\mathbf{x}}}^{(k)},\tilde{\boldsymbol{\lambda}}^{(k)})\rightarrow(\tilde{{\mathbf{x}}}^{*},\tilde{\boldsymbol{\lambda}}^{*})$ as $k\rightarrow\infty$. In addition, by the continuity of $\nabla f(\cdot)$ and $\nabla{\mathbf{g}}(\cdot)$ and the outer semi-continuity of $\mathcal{N}_{\mathcal{X}}(\cdot)$, it follows from (52) that
$\text{dist}^{2}\big{(}-\nabla f(\tilde{{\mathbf{x}}}^{*})-\nabla{\mathbf{g}}(\tilde{{\mathbf{x}}}^{*})\tilde{\boldsymbol{\lambda}}^{*},\mathcal{N}_{\mathcal{X}}(\tilde{{\mathbf{x}}}^{*})\big{)}+\text{dist}^{2}\big{(}{\mathbf{g}}(\tilde{{\mathbf{x}}}^{*}),\mathcal{N}_{\mathbb{R}_{+}^{m}}(\tilde{\boldsymbol{\lambda}}^{*})\big{)}=0,$
which easily implies that $\tilde{{\mathbf{x}}}^{*}\in\mathcal{X}^{*}$. Also, by the continuity of ${\mathbf{A}}(\cdot)-{\mathbf{b}}$, if $[{\mathbf{A}}\tilde{{\mathbf{x}}}^{*}-{\mathbf{b}}]_{j}<0$, then $[{\mathbf{A}}\tilde{{\mathbf{x}}}^{(k)}-{\mathbf{b}}]_{j}<0$ as well when $k$ is large enough, and thus $J_{A}(\tilde{\mathbf{x}}^{(k)})\subset J_{A}(\tilde{\mathbf{x}}^{*})$. Moreover, let $\tilde{I}_{g}^{(k)}=\text{supp}(\tilde{\boldsymbol{\lambda}}^{(k)})$. Then
$\textstyle\sum_{i\in\tilde{I}_{g}^{(k)}}|g_{i}(\tilde{{\mathbf{x}}}^{(k)})|^{2}+\sum_{i\in[m]\backslash\tilde{I}_{g}^{(k)}}\big{(}[g_{i}(\tilde{{\mathbf{x}}}^{(k)})]_{+}\big{)}^{2}=\text{dist}^{2}\big{(}{\mathbf{g}}(\tilde{{\mathbf{x}}}^{(k)}),\mathcal{N}_{\mathbb{R}_{+}^{m}}(\tilde{\boldsymbol{\lambda}}^{(k)})\big{)}\leq 1/k.$
Now let $k$ be large enough such that for any $i\not\in J_{g}(\tilde{{\mathbf{x}}}^{*})$, i.e., $g_{i}(\tilde{{\mathbf{x}}}^{*})<0$, it holds that $g_{i}(\tilde{{\mathbf{x}}}^{(k)})\leq\frac{1}{2}g_{i}(\tilde{{\mathbf{x}}}^{*})$ and $|g_{i}(\tilde{{\mathbf{x}}}^{*})|^{2}>\frac{4}{k}$. This observation together with the above inequality implies $\text{supp}(\tilde{\boldsymbol{\lambda}}^{(k)})=\tilde{I}_{g}^{(k)}\subset J_{g}(\tilde{{\mathbf{x}}}^{*})$. Hence, we find $\tilde{{\mathbf{x}}}^{*}\in\mathcal{X}^{*}$ such that $\|\tilde{\mathbf{x}}^{(k)}-\tilde{\mathbf{x}}^{*}\|\leq\delta^{\prime}$ and $\textup{supp}(\tilde{\boldsymbol{\lambda}}^{(k)})\subset J_{g}(\tilde{\mathbf{x}}^{*})\text{ and }J_{A}(\tilde{\mathbf{x}}^{(k)})\subset J_{A}(\tilde{\mathbf{x}}^{*})$ when $k$ is large enough. This contradicts to the assumption. Therefore, the conclusion holds. ∎
Equipped with Lemma 5.6, we are able to upper bound $\|\tilde{{\mathbf{x}}}^{(t+1)}-{\mathbf{x}}({\mathbf{z}}^{(t)})\|^{2}$ in (41) using $\text{dist}^{2}({\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)},\mathcal{N}_{\mathbb{R}_{+}^{m}}(\boldsymbol{\lambda}^{(t+1)}))$, when $\tilde{{\mathbf{x}}}^{(t+1)}$ is a near-KKT point.
@@
Suppose that Assumption 3.3 holds. Under the same conditions as in Proposition 5.5, if
@@
$\textup{dist}^{2}\big{(}-\nabla f(\tilde{{\mathbf{x}}}^{(t+1)})-\nabla{\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)})\boldsymbol{\lambda}^{(t+1)},\mathcal{N}_{\mathcal{X}}(\tilde{{\mathbf{x}}}^{(t+1)})\big{)}$ (53)
@@
$+\;\textup{dist}^{2}\big{(}{\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)}),\mathcal{N}_{\mathbb{R}_{+}^{m}}(\boldsymbol{\lambda}^{(t+1)})\big{)}$ $\leq R^{2}(\delta),$
where $\delta$ is from Assumption 3.3 and $R(\cdot)$ is from Lemma 5.6, then
@@
$\|{\mathbf{x}}({\mathbf{z}}^{(t)})-\tilde{{\mathbf{x}}}^{(t+1)}\|^{2}\leq$ $\leavevmode\nobreak\ \frac{\gamma K}{p-L}\textup{dist}^{2}\big{(}{\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)},\mathcal{N}_{\mathbb{R}_{+}^{m}}(\boldsymbol{\lambda}^{(t+1)})\big{)}.$ (54)
Proof.
From (17), we have $\tilde{\boldsymbol{\xi}}^{(t+1)}:=-\nabla_{\mathbf{x}}\mathcal{L}_{p}(\tilde{{\mathbf{x}}}^{(t+1)},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)})\in\mathcal{N}_{\mathcal{X}}(\tilde{{\mathbf{x}}}^{(t+1)}).$ Since $\mathcal{L}_{p}$ is $K$-smooth in ${\mathbf{x}}$, for any ${\mathbf{x}}\in\mathcal{X}$, we have
$\leavevmode\nobreak\ \mathcal{L}_{p}({\mathbf{x}},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)})-\mathcal{L}_{p}(\tilde{{\mathbf{x}}}^{(t+1)},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)})$
$\leq$ $-\,\langle\tilde{\boldsymbol{\xi}}^{(t+1)},{\mathbf{x}}-\tilde{{\mathbf{x}}}^{(t+1)}\rangle+\frac{K}{2}\|{\mathbf{x}}-\tilde{{\mathbf{x}}}^{(t+1)}\|^{2}.$ (55)
By the $(p-L)$-strongly convexity of $\mathcal{L}_{p}$ in ${\mathbf{x}}$, we have
Adding (55) and (56) with ${\mathbf{x}}={\mathbf{x}}({\mathbf{z}}^{(t)})$ shows that, for any ${\mathbf{x}}\in\mathcal{X}$,
$\frac{p-L}{2}\|{\mathbf{x}}({\mathbf{z}}^{(t)})-\tilde{{\mathbf{x}}}^{(t+1)}\|^{2}\leq$ $\leavevmode\nobreak\ \mathcal{L}_{p}({\mathbf{x}}({\mathbf{z}}^{(t)}),{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)})-\mathcal{L}_{p}({\mathbf{x}},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)})$
$-\langle\tilde{\boldsymbol{\xi}}^{(t+1)},{\mathbf{x}}-\tilde{{\mathbf{x}}}^{(t+1)}\rangle+\frac{K}{2}\|{\mathbf{x}}-\tilde{{\mathbf{x}}}^{(t+1)}\|^{2}.$ (57)
Recall that $\|\boldsymbol{\lambda}^{(t+1)}\|\leq M_{\boldsymbol{\lambda}}$. According to Lemma 5.6 and (53), there must exist ${\mathbf{x}}^{*}\in\mathcal{X}^{*}$ such that $\|\tilde{\mathbf{x}}^{(t+1)}-{\mathbf{x}}^{*}\|\leq\delta$, $I_{g}^{(t+1)}=\textup{supp}(\boldsymbol{\lambda}^{(t+1)})\subset J_{g}({\mathbf{x}}^{*})$ and $I_{A}^{(t+1)}=J_{A}(\tilde{\mathbf{x}}^{(t+1)})\subset J_{A}({\mathbf{x}}^{*})$. Let $\bar{{\mathbf{x}}}^{(t+1)}$ be the projection of $\tilde{\mathbf{x}}^{(t+1)}$ onto $\mathcal{S}(I_{g}^{(t+1)},I_{A}^{(t+1)})$, so $\textup{dist}(\tilde{\mathbf{x}}^{(t+1)},\mathcal{S}(I_{g}^{(t+1)},I_{A}^{(t+1)}))=\|\tilde{\mathbf{x}}^{(t+1)}-\bar{{\mathbf{x}}}^{(t+1)}\|$. As a result, it holds that
$\leavevmode\nobreak\ g_{i}(\bar{{\mathbf{x}}}^{(t+1)})=0,\;i\in I_{g}^{(t+1)},\;g_{i}(\bar{{\mathbf{x}}}^{(t+1)})\leq 0,\;i\in[m]\backslash I_{g}^{(t+1)},$ (58)
$\leavevmode\nobreak\ [{\mathbf{A}}\bar{{\mathbf{x}}}^{(t+1)}-{\mathbf{b}}]_{j}=0,\;j\in I_{A}^{(t+1)},\;[{\mathbf{A}}\bar{{\mathbf{x}}}^{(t+1)}-{\mathbf{b}}]_{j}\leq 0,\;j\in[l]\backslash I_{A}^{(t+1)},$ (59)
By Assumption 3.3 and the fact that $I_{A}^{(t+1)}=J_{A}(\tilde{\mathbf{x}}^{(t+1)})$, we have
$\|\bar{{\mathbf{x}}}^{(t+1)}-\tilde{{\mathbf{x}}}^{(t+1)}\|^{2}\leq\gamma\cdot\text{dist}^{2}\big{(}{\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)}),\mathcal{N}_{\mathbb{R}_{+}^{m}}(\boldsymbol{\lambda}^{(t+1)})\big{)}.$ (60)
$\frac{p-L}{2}\|{\mathbf{x}}({\mathbf{z}}^{(t)})-\tilde{{\mathbf{x}}}^{(t+1)}\|^{2}\leq$ $\leavevmode\nobreak\ \mathcal{L}_{p}({\mathbf{x}}({\mathbf{z}}^{(t)}),{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)})-\mathcal{L}_{p}(\bar{{\mathbf{x}}}^{(t+1)},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)})$
$+\frac{\gamma K}{2}\text{dist}^{2}\big{(}{\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)}),\mathcal{N}_{\mathbb{R}_{+}^{m}}(\boldsymbol{\lambda}^{(t+1)})\big{)}.$ (62)
$\leavevmode\nobreak\ \mathcal{L}_{p}({\mathbf{x}}({\mathbf{z}}^{(t)}),{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)})-\mathcal{L}_{p}(\bar{{\mathbf{x}}}^{(t+1)},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)})$
where the last inequality is because $\bar{{\mathbf{x}}}^{(t+1)}$ is a feasible solution to the minimization problem in (14) while ${\mathbf{x}}({\mathbf{z}}^{(t)})$ is its optimal solution. Applying (63) to (62) gives the desired result. ∎
By Proposition 5.7, we can bound the negative term $-6p\theta\|\tilde{{\mathbf{x}}}^{(t+1)}-{\mathbf{x}}({\mathbf{z}}^{(t)})\|^{2}$ in (41) by $\textup{dist}^{2}({\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)},\mathcal{N}_{\mathbb{R}_{+}^{m}}(\boldsymbol{\lambda}^{(t+1)}))$ and choosing an appropriate $\theta$ when (51) holds. The following bound will be used when (51) does not hold.
@@
$\|\tilde{{\mathbf{x}}}^{(t+1)}-{\mathbf{x}}({\mathbf{z}}^{(t)})\|^{2}\leq\frac{2M_{\boldsymbol{\lambda}}}{p-L}\textup{dist}\left({\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)},\mathcal{N}_{\mathbb{R}_{+}^{m}}(\boldsymbol{\lambda}^{(t+1)})\right).$ (64)
Proof.
By the $(p-L)$-strong convexity of $\mathcal{L}_{p}$ in ${\mathbf{x}}$, we have
$\leavevmode\nobreak\ \frac{p-L}{2}\|{\mathbf{x}}({\mathbf{z}}^{(t)})-\tilde{{\mathbf{x}}}^{(t+1)}\|^{2}\leq\mathcal{L}_{p}({\mathbf{x}}({\mathbf{z}}^{(t)}),{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)})-\mathcal{L}_{p}(\tilde{{\mathbf{x}}}^{(t+1)},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)}),$
$\leavevmode\nobreak\ \frac{p-L}{2}\|\tilde{{\mathbf{x}}}^{(t+1)}-{\mathbf{x}}({\mathbf{z}}^{(t)})\|^{2}\leq\mathcal{L}_{p}(\tilde{{\mathbf{x}}}^{(t+1)},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}({\mathbf{z}}^{(t)}))-\mathcal{L}_{p}({\mathbf{x}}({\mathbf{z}}^{(t)}),{\mathbf{z}}^{(t)},\boldsymbol{\lambda}({\mathbf{z}}^{(t)})),$
where $\boldsymbol{\lambda}({\mathbf{z}}^{(t)})$ is a vector of Lagrangian multipliers corresponding to ${\mathbf{x}}({\mathbf{z}}^{(t)})$ defined in (14). Notice that by the strong duality, $\boldsymbol{\lambda}({\mathbf{z}}^{(t)})\in\mathop{\mathrm{arg\,max}{}}_{\boldsymbol{\lambda}\geq{\mathbf{0}}}\mathcal{L}_{p}({\mathbf{x}}({\mathbf{z}}^{(t)}),{\mathbf{z}}^{(t)},\boldsymbol{\lambda})$. Hence, adding the above two inequalities gives
$(p-L)\cdot\|\tilde{{\mathbf{x}}}^{(t+1)}-{\mathbf{x}}({\mathbf{z}}^{(t)})\|^{2}\leq$ $\leavevmode\nobreak\ \mathcal{L}_{p}(\tilde{{\mathbf{x}}}^{(t+1)},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}({\mathbf{z}}^{(t)}))-\mathcal{L}_{p}(\tilde{{\mathbf{x}}}^{(t+1)},{\mathbf{z}}^{(t)},\boldsymbol{\lambda}^{(t+1)})$
$=$ $\leavevmode\nobreak\ \langle\boldsymbol{\lambda}({\mathbf{z}}^{(t)})-\boldsymbol{\lambda}^{(t+1)},{\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)})\rangle$
$\leq$ $\leavevmode\nobreak\ \textstyle\sum_{i\in\mathrm{supp}(\boldsymbol{\lambda}^{(t+1)})}|\lambda_{i}({\mathbf{z}}^{(t)})-\lambda_{i}^{(t+1)}|\cdot|g_{i}(\tilde{{\mathbf{x}}}^{(t+1)})|$
$+\textstyle\sum_{i\in[m]\backslash\mathrm{supp}(\boldsymbol{\lambda}^{(t+1)})}\lambda_{i}({\mathbf{z}}^{(t)})\cdot[g_{i}(\tilde{{\mathbf{x}}}^{(t+1)})]_{+}$
$\leq$ $\leavevmode\nobreak\ \|\boldsymbol{\lambda}({\mathbf{z}}^{(t)})-\boldsymbol{\lambda}^{(t+1)}\|\cdot\text{dist}\big{(}{\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)}),\mathcal{N}_{\mathbb{R}_{+}^{m}}(\boldsymbol{\lambda}^{(t+1)})\big{)},$ (65)
where the last inequality follows from the Cauchy-Schwarz inequality.
$\textstyle\sum_{i=1}^{m}\lambda_{i}({\mathbf{z}}^{(t)})\cdot g_{i}({\mathbf{x}}_{\text{feas}})\quad\geq\;\;\;\;$ $\leavevmode\nobreak\ \textstyle\sum_{i=1}^{m}\lambda_{i}({\mathbf{z}}^{(t)})\left(g_{i}({\mathbf{x}}({\mathbf{z}}^{(t)}))+\big{\langle}\nabla g_{i}({\mathbf{x}}({\mathbf{z}}^{(t)})),{\mathbf{x}}_{\text{feas}}-{\mathbf{x}}({\mathbf{z}}^{(t)})\big{\rangle}\right)$
[…]
$\leavevmode\nobreak\ \textstyle\|\boldsymbol{\lambda}({\mathbf{z}}^{(t)})\|_{1}\cdot\min_{i\in[m]}[-g_{i}({\mathbf{x}}_{\text{feas}})]$
$\leq$ $\leavevmode\nobreak\ \langle\nabla f({\mathbf{x}}({\mathbf{z}}^{(t)}))+p({\mathbf{x}}({\mathbf{z}}^{(t)})-{\mathbf{z}}^{(t)}),{\mathbf{x}}_{\text{feas}}-{\mathbf{x}}({\mathbf{z}}^{(t)})\rangle$
$\leq$ $\leavevmode\nobreak\ \big{(}\|\nabla f({\mathbf{x}}({\mathbf{z}}^{(t)}))\|+p\cdot\|{\mathbf{x}}({\mathbf{z}}^{(t)})-{\mathbf{z}}^{(t)}\|\big{)}\cdot\|{\mathbf{x}}_{\text{feas}}-{\mathbf{x}}({\mathbf{z}}^{(t)})\|\leq(B_{f}+pD_{\mathcal{X}})\cdot D_{\mathcal{X}},$
The following result refines (41) using (54).

### v2

@@ 5 Convergence Analysis for the iMELa Method
$\tau_{t}=\tau=\frac{p-L}{4B_{g}^{2}}\text{ and }\theta_{t}=\theta\leq\frac{p-L}{18p},\penalty 10000\ \forall\,t\geq 0.$
@@
$\phi^{t}-\phi^{t+1}\geq$ $\penalty 10000\ \frac{p-L}{4}\|{\mathbf{x}}^{(t)}-\tilde{{\mathbf{x}}}^{(t+1)}\|^{2}-6p\theta\|\tilde{{\mathbf{x}}}^{(t+1)}-{\mathbf{x}}({\mathbf{z}}^{(t)})\|^{2}$
@@
$+\frac{p}{6\theta}\|{\mathbf{z}}^{(t)}-{\mathbf{z}}^{(t+1)}\|^{2}+\tau\,\textup{dist}\big({\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)}),\mathcal{N}_{\mathbb{R}_{+}^{m}}(\boldsymbol{\lambda}^{(t+1)})\big)^{2}-\frac{\epsilon_{t}^{2}}{2(p-L)}.$ (30)
According to Proposition 5.2, to ensure a sufficient decrease of $\phi_{t}$, it is necessary to control the negative term $6p\theta\|\tilde{{\mathbf{x}}}^{(t+1)}-{\mathbf{x}}({\mathbf{z}}^{(t)})\|^{2}$ appearing in (30). To this end, we employ the local EBC with exponent $d$ in Assumption 3.3. Specifically, we show in the following proposition that, when $\tilde{{\mathbf{x}}}^{(t+1)}$ is a near-KKT point, the local EBC allows us to upper bound $\|\tilde{{\mathbf{x}}}^{(t+1)}-{\mathbf{x}}({\mathbf{z}}^{(t)})\|^{2}$ using a multiple of $\text{dist}({\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)}),\mathcal{N}_{\mathbb{R}_{+}^{m}}(\boldsymbol{\lambda}^{(t+1)}))^{2/d}$. The proof is presented in Section C.1.
@@
Suppose that Assumption 3.3 holds. There exists a constant $R(\delta)$, depending only on the constant $\delta$ in Assumption 3.3, such that, under the same conditions as in Proposition 5.2, if
@@
$\textup{dist}\big(-\nabla f(\tilde{{\mathbf{x}}}^{(t+1)})-\nabla{\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)})\boldsymbol{\lambda}^{(t+1)},\mathcal{N}_{\mathcal{X}}(\tilde{{\mathbf{x}}}^{(t+1)})\big)^{2}$
@@
$\quad+\penalty 10000\ \textup{dist}\big({\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)}),\mathcal{N}_{\mathbb{R}_{+}^{m}}(\boldsymbol{\lambda}^{(t+1)})\big)^{2}\leq R^{2}(\delta),$ (31)
it holds that
@@
$\|{\mathbf{x}}({\mathbf{z}}^{(t)})-\tilde{{\mathbf{x}}}^{(t+1)}\|^{2}\leq$ $\penalty 10000\ \frac{\gamma^{1/d}K}{p-L}\textup{dist}\big({\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)}),\mathcal{N}_{\mathbb{R}_{+}^{m}}(\boldsymbol{\lambda}^{(t+1)})\big)^{2/d}.$ (32)
Thanks to Proposition 5.3, when the local regularity condition (31) holds, the negative term $6p\theta\|\tilde{{\mathbf{x}}}^{(t+1)}-{\mathbf{x}}({\mathbf{z}}^{(t)})\|^{2}$ in (30) can be bounded above by a multiple of the quantity $\text{dist}({\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)}),\mathcal{N}_{\mathbb{R}_{+}^{m}}(\boldsymbol{\lambda}^{(t+1)}))^{2/d}$ as long as $\theta_{t}=\theta$ is chosen sufficiently small. This bound follows directly from the local EBC with exponent $d$ in Assumption 3.3.
However, when (31) does not hold, the local EBC is unavailable. In this case, we resort to the following global result, which holds without any error bound assumption. See Section C.2 for its proof.
@@
$\|\tilde{{\mathbf{x}}}^{(t+1)}-{\mathbf{x}}({\mathbf{z}}^{(t)})\|^{2}\leq\frac{2M_{\boldsymbol{\lambda}}}{p-L}\textup{dist}\big({\mathbf{g}}(\tilde{{\mathbf{x}}}^{(t+1)}),\mathcal{N}_{\mathbb{R}_{+}^{m}}(\boldsymbol{\lambda}^{(t+1)})\big).$ (33)
The result of Lemma 5.4 can be viewed as a global analogue of Assumption 3.3 with exponent $d=2$, and serves as a fallback estimate when condition (31) fails. Using Proposition 5.3 and Lemma 5.4, we can ensure a sufficient decrease of $\phi_{t}$ by choosing $\theta_{t}=\theta$ and $\tau_{t}$ in Algorithm 1 carefully. This result is presented below and is the key to establish the total complexity. The proof is given in Section C.3.